# SLIFER: Investigating Performance and Robustness of Malware Detection Pipelines

### v1

@@ Abstract.
As a result of decades of research, Windows malware detection is approached through a plethora of techniques. However, there is an ongoing mismatch between academia – which pursues an optimal performances in terms of detection rate and low false alarms – and the requirements of real-world scenarios. In particular, academia focuses on combining static and dynamic analysis within a single or ensemble of models, falling into several pitfalls like (i) firing dynamic analysis without considering the computational burden it requires; (ii) discarding impossible-to-analyse samples; and (iii) analysing robustness against adversarial attacks without considering that malware detectors are complemented with more non-machine-learning components. Thus, in this paper we propose SLIFER, a novel Windows malware detection pipeline sequentially leveraging both static and dynamic analysis, interrupting computations as soon as one module triggers an alarm, requiring dynamic analysis only when needed. Contrary to the state of the art, we investigate how to deal with samples resistance to analysis, showing how much they impact performances, concluding that it is better to flag them as legitimate to not drastically increase false alarms. Lastly, we perform a robustness evaluation of SLIFER leveraging content-injections attacks, and we show that, counter-intuitively, attacks are blocked more by YARA rules than dynamic analysis due to byte artifacts created while optimizing the adversarial strategy.
Malware analysis, machine learning, pipeline, robustness
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, U.S.A.; October 14 – 18, 2024; Salt Lake City, U.S.A.isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Due to the rapid evolution of threats and skills of malware developers, the detection of Windows malware is an on-going challenge that has kept revolutionizing itself for more than two decades. To better understand the scale of this never-ending arms race, every week $\sim$ 7M  [^1] https://virustotal.com/statistics Windows malicious programs are uploaded on cloud-based antivirus engines to be analysed. To exacerbate the issue, malware developers create several variants of the same malicious program to avoid detection by signature-based antivirus, which spot threats through unique indicators observed in the past and matched in analysed samples. Thus, modern detection engines fully embrace the machine learning (ML) paradigm, by directly understanding what makes a program malicious from data, being also able to generalize across variants.
@@
While we can only glimpse the architectures developed by industrial companies through minimally-detailed white papers (Kaspersky, 2021; Saxe and Berlin, 2015; TECHNOLOGY, [n. d.]; Avira, 2017; Microsoft, [n. d.]), academic research mostly focuses on creating models with the best trade-off between detections and false alarms (Raff et al., 2018; Trizna et al., 2023; Anderson and Roth, 2018; Jindal et al., 2019; Raff et al., 2020; Gibert et al., 2020), with a strong focus on the latter they are extremely costly to handle (Hershberger, 2023; Kubovič, 2017). To do so, state-of-the-art techniques focus on single or ensemble models that separately or jointly leverage static and dynamic analysis. While the first one infer maliciousness from both the structure and the content of input samples, the second one requires programs to be executed or emulated inside an isolated environment.
@@
However, dynamic analysis is costly, requiring a clean environment at each analysis where samples can be “detonated” to manifest their behavior. Every ML architecture that uses this kind of analysis must face this mandatory cost, and also plenty of proposed approaches consider features extracted with dynamic analysis fused or stacked together (Dambra et al., 2023; Trizna, 2022) with the static ones. This not only expands the complexity of the feature extraction phase, making it more costly and time-consuming, but also the overall predictive capability might not even benefit from this addition: Dambra et al. (Dambra et al., 2023) highlight a worrisome trend where dynamic analysis is neither better or complementary to the static one, diminishing the belief of tracing execution as the Swiss knife against malware. On the other hand, most of the proposed techniques hide the crashes encountered by static analysis while extracting features, resolving the issue by discarding those impossible-to-analyse samples. While this is not a problem at the research stage, such a behavior is not admissible in production environments where an answer must always be given to the users who requested an analysis.
@@
Hence, in this work we propose SLIFER, a novel Windows malware detector that matches the needs of production environments. SLIFER is built on three components: (i) pattern-matching with YARA rules to rapidly filter out known samples; (ii) static malware detection with state-of-the-art models to capture most of the threats; and (iii) dynamic analysis with emulation to fine-tune results on more difficult samples. Input programs traverse the pipeline sequentially, halting the process at the first detection. This step reduces the needs for firing dynamic analysis, focusing its resource-demanding feature extraction process only when really needed. Static analysis is achieved with a mixture of models, combining end-to-end –less predictive, but it can not crash at inference time due to the absence of feature extraction– and feature-extraction-based –more predictive, but can crash– detection to reduce the number of crashes. Through SLIFER, which mimics an industrial malware detector, we want to answer three research questions:
@@
RQ2: which are the performances of sequential analysis, compared to the proposed single and hybrid approaches of the literature?
RQ3: what is the robustness of a pipeline of components?
RQ4: what is the overhead brought by a sequential pipeline in analysing input samples?
@@
Take-home message 1. Counter-intuitively, samples causing errors during pre-processing should advance in the pipeline, being labelled as benign in case no modules are able to analyse them. This reduces drastically the number of false alarms, and partially the detection, still being acceptable in production environments.
@@
Take-home message 3. SLIFER is more robust against static adversarial EXEmples, but not thanks to dynamic analysis. In fact, we show that artifacts generated by attacks are detected by YARA rules that were not matching the unperturbed program. Also, we show that content injection attacks might have a marginal effect on functionality, since content is displaced and retrieved at different addresses at runtime, causing the dynamic-based ML model to reduce its score.
Take-home message 4. SLIFER does not negatively impact the time needed for analysing samples, since most of the malicious programs are rapidly recognized by static analysis, requiring the dynamic one only for benignware.
The rest of the paper is organised as follows: we firstly introduce the background concepts needed to understand our work (Sect. 2), followed by the implementation details of SLIFER (Sect. 3). We then continue by describing our experiments (Sect. 4), and which results we derive from them, along with the answers to our research questions (Sect. 5). We conclude the paper by detailing the limitations of our approach (Sect. 7), possible future research directions, and final remarks (Sect. 8).
2. Background
@@
Figure 1. Graphical representation of the Windows PE format.
DOS Header + Stub. These chunks of bytes represent a valid DOS program, kept for retrocompatibility. It is useless in modern OSes.
@@
Sections. These contain the code and asset of the program to load. Usually, the first section contains the machine instructions of the software, while the others are used as storage for initialized variables, resources, and other relevant information.
Windows Malware Detection. To stop the spreading of malware, various techniques have been developed leveraging either static, dynamic, or both types of analyses.
Static analysis. This type of analysis is based on the extraction of relevant metrics from the structure and content of analysed samples (Raff et al., 2018, 2020; Anderson and Roth, 2018), without the need for executing it. The most naïve static analysis methodology is posed by pattern-matching with well-known signatures (Yara-Rules, 2019; The FLARE Team, 2020), but such a technique is not robust against all the malware variants that are released on a daily basis. Thus, ML models are trained on static features to gain general performance also on unseen samples.
Dynamic analysis. This type of analysis concentrates on characterizing the behavior of programs (Trizna et al., 2023; Jindal et al., 2019) by recording their trace of execution (spawned processes, API calls, reached websites, accessed registry keys) while being detonated inside an isolated environment, such as a sandbox, or through emulation. After having collected all the events, these are pre-processed to be fed to machine learning models as training data.
@@
Adversarial EXEmples. The rise of ML in malware detection brought in parallel the rise of Adversarial Machine Learning(Yuan et al., 2019; Wiyatno et al., 2019; Biggio et al., 2013; Biggio and Roli, 2018), which wants to exploit vulnerabilities of models creating adversarial EXEmples (Demetrio et al., 2021b, a). These are well-crafted programs tailored to fool ML malware detectors, causing impairments on end users’ devices. While limited research focuses on attacking dynamic detectors (Rosenberg et al., 2018), most of the efforts have been focused on evading static classifiers (Demetrio et al., 2021b, a; Anderson et al., 2017; Lucas et al., 2021). These attacks work by either replacing existing content or by injecting new one that disrupt the pattern learned at training time.
3. SLIFER: sequential pipeline for Windows Malware Detection
Figure 2. The architecture of SLIFER. Input traverses the pipeline by triggering each module sequentially, and computations are halted when one module raises an alert. In presence of pre-processing errors, SLIFER forwards the input to the next module. This can happen in non-end-to-end modules such as GBDT-EMBER and Nebula.
We now describe SLIFER, a Windows malware detector built on different components, leveraging both static and dynamic analysis as depicted in Fig. 2. Differently from state-of-the-art techniques that use single or ensemble models, SLIFER performs predictions by sequentially testing different modules, halting computations for a specific sample when it is detected as malicious by one component. Instead of merging together static and dynamic analysis that would require detonating samples in sandboxes, SLIFER leverages emulation to retrieve execution traces, and such operation is only required for programs that are flagged as non-malicious by all the previous modules. In this way, we minimize the time of detection for malicious files, also thanks to the order of the modules, sorted from the fastest to the slowest. When one of the modules fails to analyse a sample due to errors or crashes of pre-processing, SLIFER continues the analysis by passing the input to the next one. If no module is able to successfully compute a prediction, we flag the sample as benign. We will later show that this counter-intuitive choice is applied to keep the number of false alarms extremely low, while not reducing much the predictions on malicious samples (Sect. 5.1). We now detail all the components of SLIFER, by discussing their implementation and design choices.
3.1. Pattern-matching Detection with Signatures
@@
3.2. Machine Learning Static Analysis
@@
MalConv. Proposed by Raff et al. (Raff et al., 2018), this model has been developed to learn maliciousness directly from bytes, by taking in input whole executable files and returning probability scores. It is implemented as a convolutional neural network (CNN), starting from an embedding layer that encodes bytes inside a space where distances are defined. Input is then processed through a gated convolution layer, a global max-pooling layer, and a fully connected layer that computes the final prediction. MalConv is trained on the state-of-the-art dataset EMBER (Anderson and Roth, 2018), and it takes in input the first 1 MB of each sample. In case input programs are shorter than this amount, they are padded with a special value.
@@
GBDT-EMBER. Proposed by Anderson et al. (Anderson and Roth, 2018), this model leverages a Gradient Boosted Decision Tree (GBDT) trained on the state-of-the-art dataset EMBER (Anderson and Roth, 2018), containing features extracted from Windows Portable Executables (PEs). Features consist of eight groups, including features extracted after parsing the PE, and ”format-agnostic” features, containing characteristics obtained without the parsing. The parsed features groups are: (i) general file information, including the information obtained from the PE header; (ii) header information taken from the headers, like the target machine, the target subsystem and so on; (iii) imported functions, reporting the imported functions by library; (iv) exported functions as a list; (v) section information, comprehending the properties of each section. The format-agnostic features are a (i) byte histogram, representing the normalized counts of each byte value within the file, a (ii) byte-entropy histogram which accounts for the entropy of the byte distribution of the file, applying a sliding window over the binary and (iii) string information taken from the printable strings inside the PE.
@@
Each input sample is first analysed by MalConv, and, if the model is not rising any alert, it is passed to GBDT-EMBER. In this way, we first analyse samples with a faster model with no pre-processing, thus reducing the overall number of possible errors.
3.3. Machine Learning Dynamic Analysis
@@
The last module of SLIFER leverages dynamic analysis to trace the execution of samples, and it should help SLIFER to recognize obfuscated and packed samples that evaded the previous modules. Among the recently released models for dynamic analysis, we select Nebula (Trizna et al., 2023), a pipeline influenced by advances of Large Language Models (LLM) (Radford et al., 2018) employing the self-attention neural mechanism to analyze dynamic analysis reports and classify samples. Nebula employs Windows kernel emulation as a compromise between computational complexity and coverage. While emulation is cheaper than system virtualization, it is more prone to dynamic analysis errors. Nebula introduces domain-knowledge-influenced filters to distill behavioral reports from redundant or irrelevant information like memory addresses or PE file segment hashes. Further, the behavioral report is processed by the Transformer encoder neural network, producing a probability of maliciousness. We use pre-trained objects released by Nebula, trained on a public dataset (Trizna, 2022), comprising $\approx$75k samples collected in January 2022, spanning across seven malware types and $\approx$25k benignware samples.
3.4. SLIFER Implementation

### v3

@@ Abstract
As a result of decades of research, Windows malware detection is approached through a plethora of techniques. However, there is an ongoing mismatch between academia – which pursues an optimal performances in terms of detection rate and low false alarms – and the requirements of real-world scenarios. In particular, academia focuses on combining static and dynamic analysis within a single or ensemble of models, falling into several pitfalls like (i) firing dynamic analysis without considering the computational burden it requires; (ii) discarding impossible-to-analyze samples; and (iii) analyzing robustness against adversarial attacks without considering that malware detectors are complemented with more non-machine-learning components. Thus, in this paper we bridge these gaps, by investigating the properties of malware detectors built with multiple and different types of analysis. To do so, we develop SLIFER, a Windows malware detection pipeline sequentially leveraging both static and dynamic analysis, interrupting computations as soon as one module triggers an alarm, requiring dynamic analysis only when needed. Contrary to the state of the art, we investigate how to deal with samples that impede analyzes, showing how much they impact performances, concluding that it is better to flag them as legitimate to not drastically increase false alarms. Lastly, we perform a robustness evaluation of SLIFER. Counter-intuitively, the injection of new content is either blocked more by signatures than dynamic analysis, due to byte artifacts created by the attack, or it is able to avoid detection from signatures, as they rely on constraints on file size disrupted by attacks. As far as we know, we are the first to investigate the properties of sequential malware detectors, shedding light on their behavior in real production environment.
keywords:
[…]
1 Introduction
@@
Due to the rapid evolution of threats and skills of malware developers, the detection of Windows malware is an on-going challenge that has kept revolutionizing itself for more than two decades. To better understand the scale of this never-ending arms race, every week $\sim$ 7M [^1] https://virustotal.com/statistics Windows malicious programs are uploaded on cloud-based antivirus engines to be analyzed. To exacerbate the issue, malware developers create several variants of the same malicious program to avoid detection by signature-based antivirus, which spot threats through unique indicators observed in the past and matched in analyzed samples. Thus, modern detection engines fully embrace the machine learning (ML) paradigm, by directly inferring the maliciousness of program from data, being also able to generalize across variants.
@@
While we can only glimpse the architectures developed by industrial companies through minimally-detailed white papers (Kaspersky, 2021; Saxe and Berlin, 2015; ESET Technology, 2017; Avira, 2017; Microsoft, 2017), academic research mostly focuses on creating models with the best trade-off between detections and false alarms (Raff et al., 2018; Trizna et al., 2024; Anderson and Roth, 2018; Jindal et al., 2019; Raff et al., 2020; Gibert et al., 2020), with a strong focus on the latter they are extremely costly to handle (Hershberger, 2023; Kubovič, 2017). To do so, state-of-the-art techniques focus on single or ensemble models that separately or jointly leverage static and dynamic analysis. While the first one infer maliciousness from both the structure and the content of input samples, the second one requires programs to be executed or emulated inside an isolated environment.
@@
However, dynamic analysis is costly, requiring a clean environment at each analysis where samples can be “detonated” to manifest their behavior. Every ML architecture that uses this kind of analysis must face this mandatory cost, and also plenty of proposed approaches consider features extracted with dynamic analysis fused or stacked together (Dambra et al., 2023; Trizna, 2022) with the static ones. This not only expands the complexity of the feature extraction phase, making it more costly and time-consuming, but also the overall predictive capability might not even benefit from this addition: Dambra et al. (2023) highlight a worrisome trend where dynamic analysis is neither better or complementary to the static one, diminishing the belief of tracing execution as the Swiss knife against malware. On the other hand, most of the proposed techniques hide the crashes encountered by static analysis while extracting features, resolving the issue by discarding those impossible-to-analyze samples. While this is not a problem at the research stage, such a behavior is not admissible in production environments where an answer must always be given to the users who requested an analysis.
@@
Hence, in this work we conduct a detailed investigation on the behavior of sequential malware detection pipeline, by addressing the highlighted roadblocks of the state of the art. To do so, we leverage existing technology to build SLIFER, a malware detector that matches the needs of production environment.
SLIFER is built on three components: (i) pattern-matching with YARA rules to rapidly filter out known samples; (ii) static malware detection with state-of-the-art models to capture most of the threats; and (iii) dynamic analysis with emulation to fine-tune results on more difficult samples. Input programs traverse the pipeline sequentially, halting the process at the first detection. This step reduces the need for firing dynamic analysis, focusing its resource-demanding feature extraction process only when needed. Static analysis is achieved with a mixture of models, combining end-to-end –less predictive, but resistant to crashes due to the absence of feature extraction– and feature-extraction-based –more predictive, but likely to encounter pre-processing errors– detection to reduce the number of crashes. Through SLIFER, which mimics an industrial malware detector, we want to answer three research questions:
@@
RQ2: which are the performances of sequential analysis, compared to existing single and hybrid approaches?
RQ3: is it possible to tune the detection threshold of each component of SLIFER w.r.t. to a single validation set?
RQ4: how much detection performance and false positive rates varies when deploying calibrated pipeline w.r.t. the naïve combination of pre-trained models?
RQ5: what is the robustness of an end-to-end pipeline of pre-built components?
RQ6: what is the overhead brought by a sequential pipeline in analysing input samples?
@@
Take-home message 1. Counter-intuitively, samples causing errors during pre-processing should advance in the pipeline, being labeled as benign in case no modules are able to analyze them. This reduces drastically the number of false alarms, and partially the detection, still being acceptable in production environments.
@@
Take-home message 3. We show how to tune the detection threshold of each component of SLIFER by applying grid search, highlighting their substantial change w.r.t. to the original ones computed at their training time.
Take-home message 4. Counter-intuitively, calibrated thresholds of the components of SLIFER does not bring relevant improvements to accuracy and false positive rate.
Take-home message 5. SLIFER is more robust against static adversarial EXEmples, but not thanks to dynamic analysis. We show that artifacts generated by attacks are detected by YARA rules that were not matching the unperturbed program. We show that content injection attacks might have a marginal effect on functionality, since content is displaced and retrieved at different addresses at runtime, causing the dynamic-based ML model to reduce its score. Lastly, we highlight that some signatures, relying on constraints on the filesize, are bypassed by adversarial attacks.
Take-home message 6. SLIFER does not negatively impact the time needed for analysing samples, since most of the malicious programs are rapidly recognized by static analysis, requiring the dynamic one only for benignware.
As far as we know, we are the first to investigate the behavior of sequential pipelines like SLIFER, shortening the gap between real production environment and academia. In particular, we are the first to perform a realistic testing, not only in regular industrial environment where pre-processing errors must be considered, but also in presence of skillful attackers that try to bypass all the layers of detection.
The rest of the paper is organized as follows: we firstly introduce the background concepts needed to understand our work (Sect. 2), followed by the implementation details of SLIFER (Sect. 3). We then continue by describing our experiments (Sect. 4), and which results we derive from them, along with the answers to our research questions (Sect. 5). We conclude the paper by detailing the limitations of our approach (Sect. 7), possible future research directions, and final remarks (Sect. 8).
2 Background
@@
Figure 1: Graphical representation of the Windows PE format.
DOS Header + Stub. These unused chunks of bytes represent a valid DOS program, kept for retrocompatibility.
@@
Sections. These contain code and assets of the program to load. Usually, the first section contains machine instructions of the software, followed by storages for initialized variables, resources, and other relevant information.
Windows Malware Detection. To stop the spreading of malware, various techniques have been developed leveraging either static, dynamic, or both types of analysis.
Static analysis. This type of analysis is based on the extraction of relevant metrics from the structure and content of analyzed samples (Raff et al., 2018, 2020; Anderson and Roth, 2018), without the need for executing it. The most naïve static analysis methodology is posed by pattern-matching with well-known signatures (Yara-Rules, 2019; The FLARE Team, 2020), but such a technique is not robust against all the malware variants that are released on a daily basis. Thus, ML models are trained on static features to gain general performance also on unseen samples.
Dynamic analysis. This type of analysis concentrates on characterizing the behavior of programs (Trizna et al., 2024; Jindal et al., 2019) by recording their trace of execution (spawned processes, API calls, reached websites, accessed registry keys) while being detonated inside an isolated environment, such as a sandbox, or through emulation. After having collected all the events, these are pre-processed to be fed to machine learning models as training data.
@@
Adversarial EXEmples. The rise of ML in malware detection brought in parallel the rise of Adversarial Machine Learning (Yuan et al., 2019; Wiyatno et al., 2019; Biggio et al., 2013; Biggio and Roli, 2018). In particular, in the domain of malware detection, this translates in the creation of adversarial EXEmples (Demetrio et al., 2021b, a), carefully-perturbed programs tailored to fool ML-based detectors. While limited research focuses on attacking dynamic detectors (Rosenberg et al., 2018), most of the efforts have been focused on evading static classifiers (Demetrio et al., 2021b, a; Anderson et al., 2017; Lucas et al., 2021). These attacks work by either replacing existing content or by injecting new one that disrupt the pattern learned at training time.
3 SLIFER: Sequential Pipeline For Windows Malware Detection
Figure 2: The architecture of SLIFER. Input traverses the pipeline by triggering each module sequentially, and computations are halted when one module raises an alert. In presence of pre-processing errors, SLIFER forwards the input to the next module. This can happen in non-end-to-end modules such as GBDT-EMBER and Nebula.
We now describe SLIFER, a Windows malware detector built on different components, leveraging both static and dynamic analysis as depicted in Fig. 2. Different from state-of-the-art techniques that use single or ensemble models, SLIFER performs predictions by sequentially testing different modules, halting computations for a specific sample when it is detected as malicious by one component. Instead of merging together static and dynamic analysis that would require detonating samples in sandboxes, SLIFER leverages emulation to retrieve execution traces, and such operation is only required for programs that are flagged as non-malicious by all the previous modules. In this way, we minimize the time of detection for malicious files, also thanks to the order of the modules, sorted from the fastest to the slowest. When one of the modules fails to analyze a sample due to errors or crashes of pre-processing, SLIFER continues the analysis by passing the input to the next one. If no module is able to successfully compute a prediction, we flag the sample as benign. We will later show that this counter-intuitive choice is applied to keep the number of false alarms extremely low, without reducing much the predictions on malicious samples (Sect. 5.1). We now detail all the components of SLIFER, by discussing their implementation and design choices.
3.1 Pattern-matching Detection with Signatures
@@
3.2 Machine Learning Static Analysis
@@
MalConv. Proposed by Raff et al. (2018), this model has been developed to learn maliciousness directly from bytes, by taking in input whole executable files and returning probability scores. It is implemented as a convolutional neural network (CNN), starting from an embedding layer that encodes bytes inside a space where distances are defined. Input is then processed through a gated convolution layer, a global max-pooling layer, and a fully connected layer that computes the final prediction. MalConv is trained on the state-of-the-art dataset EMBER (Anderson and Roth, 2018), and it takes in input the first 1 MB of each sample. In case input programs are shorter than this amount, they are padded with a special value.
@@
GBDT-EMBER. Proposed by Anderson and Roth (2018), this model leverages a Gradient Boosted Decision Tree (GBDT) trained on the state-of-the-art dataset EMBER (Anderson and Roth, 2018), containing features extracted from Windows Portable Executables (PEs). Features consist of eight groups, including features extracted after parsing the PE, and "format-agnostic" features, containing characteristics obtained without the parsing. The parsed features groups are: (i) general file information, including the information obtained from the PE header; (ii) header information taken from the headers, like the target machine, the target subsystem and so on; (iii) imported functions, reporting the imported functions by library; (iv) exported functions as a list; (v) section information, comprehending the properties of each section. The format-agnostic features are a (i) byte histogram, representing the normalized counts of each byte value within the file, a (ii) byte-entropy histogram which accounts for the entropy of the byte distribution of the file, applying a sliding window over the binary and (iii) string information taken from the printable strings inside the PE.
@@
Each input sample is first analyzed by MalConv, and, if the model is not raising any alert, it is passed to GBDT-EMBER. In this way, we first analyze samples with a faster model with no pre-processing, thus reducing the overall number of possible errors.
3.3 Machine Learning Dynamic Analysis
@@
The last module of SLIFER leverages dynamic analysis to trace the execution of samples, and it should help SLIFER to recognize obfuscated and packed samples that evaded the previous modules. Among the recently released models for dynamic analysis, we select Nebula (Trizna et al., 2024), a pipeline influenced by advances of Large Language Models (LLM) (Radford et al., 2018) employing the self-attention neural mechanism to analyze dynamic analysis reports and classify samples. Nebula employs Windows kernel emulation as a compromise between computational complexity and coverage. While emulation is cheaper than system virtualization, it is more prone to dynamic analysis errors. Nebula introduces domain-knowledge-influenced filters to distill behavioral reports from redundant or irrelevant information like memory addresses or PE file segment hashes. Further, the behavioral report is processed by the Transformer encoder neural network, producing a probability of maliciousness. We use pre-trained objects released by Nebula, trained on a public dataset (Trizna, 2022), comprising $\approx$75k samples collected in January 2022, spanning across seven malware types and $\approx$25k benignware samples.
3.4 SLIFER Implementation